# EviNet:Towards Open-World Graph Learning via Evidential Reasoning Network

### v2

@@ 1. Introduction
Graphs capture relationships among entities in many real-world scenarios, including social networks (Khanam et al., 2023; Wang et al., 2024b), recommendation systems (Wu et al., 2022, 2025), and financial fraud detection (Wang et al., 2022, 2025). To date, researchers have developed a wealth of graph learning models, such as graph neural networks, to learn node representations on graphs (Kipf and Welling, 2017; Hamilton et al., 2017; veličković2018graph). These methods often learn the representation of a node by aggregating information from its neighbors and updating the model parameters by its label. An underlying assumption for training these graph learning models is that all possible labels are known a priori in a closed label set during training, which we term learning in a closed-world environment.
@@ 3.1. Framework Overview
An overview of our proposed framework is presented in Figure 2, which consists of two key modules: (M1) Dissonance Reasoning and (M2) Vacuity Reasoning. The Dissonance Reasoning module tackles (C1) in-distribution misclassification detection by generating
class embeddings for known (in-distribution) classes from nodes and computing a dissonance score, which indicates the conflicting evidence towards multiple known classes, using the node embeddings and class embeddings.
@@ 4.2. Effectiveness Results
| Dataset | Metric | ODIN | MaxLogit | Mahalanobis | Energy | GNNSafe | CRL | DOCTOR | GKDE | GPN | EviNet | |
| Amazon -photo | IDC | Acc () | ||||||||||
| MD | AURC () | |||||||||||
| OODD | FPR95 () | |||||||||||
| AUROC () | ||||||||||||
| Amazon -computer | IDC | Acc () | ||||||||||
| MD | AURC () | |||||||||||
| OODD | FPR95 () | |||||||||||
| AUROC () | ||||||||||||
| Coauthor -cs | IDC | Acc () | ||||||||||
| MD | AURC () | |||||||||||
| OODD | FPR95 () | |||||||||||
| AUROC () | ||||||||||||
| Coauthor -physics | IDC | Acc () | ||||||||||
| MD | AURC () | |||||||||||
| OODD | FPR95 () | |||||||||||
| AUROC () | ||||||||||||
| Wiki-cs | IDC | Acc () | ||||||||||
| MD | AURC () | |||||||||||
| OODD | FPR95 () | |||||||||||
| AUROC () | ||||||||||||
@@ Appendix A Symbols and Notations
| Dataset | Metric | ODIN | MaxLogit | Mahalanobis | Energy | GNNSafe | CRL | DOCTOR | GKDE | GPN | EviNet | |
| Amazon -photo | IDC | Acc () | ||||||||||
| MD | AURC () | |||||||||||
| AUROC () | ||||||||||||
| AUPR () | ||||||||||||
| OODD | FPR95 () | |||||||||||
| AUROC () | ||||||||||||
| AUPR () | ||||||||||||
@@ Appendix E Scalability Analysis
| Method | Training Time (s)/epoch | Inference Time (s)/epoch | Memory Usage (MB) |
| ODIN | 0.020 | 2.66 | 862 |
| MaxLogit | 0.019 | 2.4 | 862 |
| GNNSafe | 0.032 | 2.46 | 1095 |
| CRL | 0.020 | 2.5 | 863 |
| DOCTOR | 0.019 | 2.5 | 862 |
| GKDE | 21.17 | 2.4 | 878 |
| GPN | / | / | / |

### v3

@@ 1. Introduction
Graphs capture relationships among entities in many real-world scenarios, including social networks (Khanam et al., 2023; Wang et al., 2024b), recommendation systems (Wu et al., 2022, 2025), and financial fraud detection (Wang et al., 2022, 2025). To date, researchers have developed a wealth of graph learning models, such as graph neural networks, to learn node representations on graphs (Kipf and Welling, 2017; Hamilton et al., 2017; Velickovic et al., 2018). These methods often learn the representation of a node by aggregating information from its neighbors and updating the model parameters by its label. An underlying assumption for training these graph learning models is that all possible labels are known a priori in a closed label set during training, which we term learning in a closed-world environment.
@@ 3.1. Framework Overview
An overview of our proposed framework is presented in Figure 2, which consists of two key modules: (M1) Dissonance Reasoning and (M2) Vacuity Reasoning. The Dissonance Reasoning module tackles (C1) in-distribution misclassification detection by generating class embeddings for known (in-distribution) classes from nodes and computing a dissonance score, which indicates the conflicting evidence towards multiple known classes, using the node embeddings and class embeddings.
@@ 4.2. Effectiveness Results
| Dataset | Metric | ODIN | MaxLogit | Mahalanobis | Energy | GNNSafe | CRL | DOCTOR | GKDE | GPN | EviNet | |||
IDC | Acc () | ||||||||||||
| MD | AURC () | |||||||||||||
| OODD | FPR95 () | |||||||||||||
| AUROC () | ||||||||||||||
IDC | Acc () | ||||||||||||
| MD | AURC () | |||||||||||||
| OODD | FPR95 () | |||||||||||||
| AUROC () | ||||||||||||||
IDC | Acc () | ||||||||||||
| MD | AURC () | |||||||||||||
| OODD | FPR95 () | |||||||||||||
| AUROC () | ||||||||||||||
IDC | Acc () | ||||||||||||
| MD | AURC () | |||||||||||||
| OODD | FPR95 () | |||||||||||||
| AUROC () | ||||||||||||||
| Wiki-cs | IDC | Acc () | ||||||||||||
| MD | AURC () | |||||||||||||
| OODD | FPR95 () | |||||||||||||
| AUROC () | ||||||||||||||
@@ Appendix A Symbols and Notations
| Dataset | Metric | ODIN | MaxLogit | Mahalanobis | Energy | GNNSafe | CRL | DOCTOR | GKDE | GPN | EviNet | |||
IDC | Acc () | ||||||||||||
| MD | AURC () | |||||||||||||
| AUROC () | ||||||||||||||
| AUPR () | ||||||||||||||
| OODD | FPR95 () | |||||||||||||
| AUROC () | ||||||||||||||
| AUPR () | ||||||||||||||
@@ Appendix E Scalability Analysis
| Method |
| ODIN | 0.020 | 2.66 | 862 | ||||||
| MaxLogit | 0.019 | 2.4 | 862 | ||||||
| GNNSafe | 0.032 | 2.46 | 1095 | ||||||
| CRL | 0.020 | 2.5 | 863 | ||||||
| DOCTOR | 0.019 | 2.5 | 862 | ||||||
| GKDE | 21.17 | 2.4 | 878 | ||||||
| GPN | / | / | / | ||||||